OpenStudy (anonymous):

If a central angle of measure 30° is subtended by a circular arc of length 6 meters, as is illustrated below, how many meters in length is the radius of the circle? How do you do this?

OpenStudy (anonymous):

the diagram ?

OpenStudy (anonymous):

no i got it circumference is calculated using the formula \[( \theta \div 360 ) \times 2 \pi \times radius\]

OpenStudy (amistre64):

s = (a) r r = s/(a) right?

OpenStudy (anonymous):

1 / 12 of 2 pi * r = 6 1/6 * pi *r = 6 pi * r = 36 r = 36 / pi r = 11.45 ( approx - taking pi as 3.142 )

OpenStudy (amistre64):

6 ----- = 36/pi = .... yeah pi/6